OpenStudy (anonymous):

Choose the system of inequalities that matches the graph below. MULTIPLE CHOICE

OpenStudy (anonymous):

x – y ≤ 5 5x + 6y ≤ 0 x – y ≥ 5 5x + 6y ≤ 0 x – y < 5 5x + 6y < 0 x – y > 5 5x + 6y < 0

OpenStudy (goten77):

i know there is 2 possible ways to do this but i dont know the short way.. but the long way is. you get y by itself on all of theese then you imagine its y= whatever and depending on the sign you can tell if the graph is under or below it

OpenStudy (anonymous):

okay

OpenStudy (goten77):

also since the 1st 2 equasions say equal to that means that the line on the graph would include those values and be a full straight line with no dash's in it.. a dashed line means that it does not equal that value so the last 2 options of this are not possible

OpenStudy (anonymous):

so that eliminates C and D!

OpenStudy (goten77):

yes

OpenStudy (anonymous):

Judging by the intercepts, I'm pretty sure it's B. what do you think?

OpenStudy (goten77):

hmm x – y ≥ 5 can be re written as \[x-5 \ge y\]

OpenStudy (goten77):

\[y \le x-5\]

OpenStudy (goten77):

i would then say y=x-5 and that line has the y intercept at-5 and the area highlighted would always be below the curve so B is right

OpenStudy (goten77):

answer choice a would be the area shaded above y=x-5 and that is not shown in the picture

OpenStudy (anonymous):

so it must be b